Micros Unlimited corporation (a local distributor of micro computers) sells one of its 586-based system on a volume

discount as follows: 1 to 5 units at a price of $4495 per unit, units 6 through 10 at a price of $ 4095 per unit, and

any unit in excess of 10 at a price of $3775 per unit.

i. Determine the cost function and draw its graph.

ii. What are the total and average cost s of 4 units ?

iii. What are the total and average cost s of 15 units?

iv. How many units were purchased if the total charge was $58,050 ?

v. How many units were purchased if the average charge was $4035?


1
Expert's answer
2021-11-02T15:12:03-0400

i)

p - price, x - units

"p(x)=\begin{cases}
 4495 & 1\le x\le 5 \\
 4095 & 6\le x \le 10\\
3775&x>10
\end{cases}"



ii)

average cost:

"p(4)=\$4495"


total cost:

"4\cdot 4495=\$17980"


iii)

average cost:

"\frac{5\cdot4495+5\cdot4095+5\cdot3775}{15}=\$4121.67"


total cost:

"15\cdot4121.67=\$61825"


iv)

"58050=5\cdot4495+5\cdot4095+4\cdot3775"

were purchased:

"5+5+4=14" units


v)

"4035=\frac{5\cdot4495+5\cdot4095+3775n}{N}"

"N=10+n"

Then:

"4035(10+n)=42950+3775n"

"260n=2600"

"n=10"


were purchased:

"N=5+5+10=20" units